OpenStudy (anonymous):

Why can't 57 be written as the sum of two squares? I was under the impression that only numbers of the form 3 mod 4 could not be written in this way, and 57 does not equal 3 mod 4? Unless I'm missing something majorly obvious Thanks

OpenStudy (kc_kennylau):

Because 1+1=2, 1+4=5, 1+9=10, 1+16=17, 1+25=26, 1+36=37, 1+49=50, 4+4=8, 4+9=13, 4+16=20, 4+25=29, 4+36=40, 4+49=53, 9+9=18, 9+16=25, 9+25=34, 9+36=44, 16+16=32, 16+25=31, 16+36=52, 25+25=50

OpenStudy (kc_kennylau):

I don't really think that you can predict which numbers can be written as the sum of two squares...

OpenStudy (kc_kennylau):

Wait, something from Wikipedia: A positive integer can be represented as a sum of two squares precisely if its prime factorization contains no odd powers of primes of the form 4k + 3. This is generalized by Waring's problem.

OpenStudy (kc_kennylau):

19 is in the form of 4k + 3

OpenStudy (anonymous):

ahh okay, it's the prime factorisation which can't be of the form 3 mod 4. Great thanks for that!!